ROTATIONAL SPECTRA OF p-, m-, AND o-CYANOPHENOL AND INTERNAL ROTATION OF p-CYANOPHENOL

\underline{ANDREW R. CONRAD}, NATHAN Z. BAREFOOT, and MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, Ohio 44242.

Rotational spectra of p-, m-, and o-cyanophenol have been measured in the range of 10.5 - 21 GHz and fit using Watson’s A-reduction Hamiltonian coupled with nuclear quadrupole coupling interaction terms for the ^{14}N nuclei. Ab initio calculations at the MP2/6-311++G** level predict the cis conformers of m- and o-cyanophenol to be more stable due to the intramolecular hydrogen bonding interaction between the hydroxyl hydrogen and the cyano nitrogen. We recorded 14 a- and b-type rotational transitions for cis m-cyanophenol and 16 a- and b-type rotational transitions for trans mcyanophenol. The rotational constants are A = 3408.9200(2) MHz, B = 1205.8269(2) MHz, and C = 890.6672(1) MHz and A = 3403.1196(3) MHz, B = 1208.4903(2) MHz, and C = 891.7241(2) MHz for the cis and trans species respectivly. We recorded 25 a- and b-type rotational transitions for cis o-cyanophenol; the rotational constants are A = 3053.758(2) MHz, B = 1511.2760(3) MHz, and C = 1010.7989(2) MHz. The trans conformer of o-cyanophenol was not observed. Rotational transitions of the p-cyanophenol monomer are split due to the symmetric internal rotation of the hydroxyl group with respect to the aromatic ring. We recorded 25 a- and b-type rotational transitions for p-cyanophenol; the b-type transitions are split by 40 MHz. The rotational constants are A = 5612.96(2) MHz, B = 990.4283(6) MHz, and C = 841.9364(6) MHz and the ground state spitting ΔE is 20.1608(6) MHz.